OpenStudy (anonymous):

Help Please!!!!!!!!!!!! Ill Give Medal!!!!!!!! Which step should be completed first to solve the equation? 8x -12 = 13 A. Add 12 to both sides. B. Add 13 to both sides. C. Divide both sides by 8. D. Subtract 13 from both sides.

OpenStudy (anonymous):

@BunnyBree

OpenStudy (anonymous):

The first step would be to add 12 to both sides. Adding 12 to both sides would isolate the variable on one side, allowing you to divide and solve the equation. :D

OpenStudy (anonymous):

:D thanks can you help me with some more????

OpenStudy (anonymous):

Sure. (:

OpenStudy (anonymous):

What is the solution to the equation \[8 - 2b = -\frac{ 2 }{ 3 }(12b + 6)\] B = ???

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

The first thing we need to do is distribute -2/3 to everything in the parenthesis. -2/3 * 12b = -8b -2/3 * 6 = -4 Now we have: 8 - 2b = -8b - 4 Add 2b to both sides. 8 = -6b - 4 Add 4 to both sides. 12 = -6b Divide each side by -6. b = -2 <--- Answer. (: Check: 8-2b=-2/3(12b+6) 8 - 2(-2) = -2/3(12(-2) + 6) 8 + 4 = -2/3(-24 + 6) 12 = -2/3 (-18) 12 = 12

OpenStudy (anonymous):

:D yay thanks again hah i have like 3 more :)

OpenStudy (anonymous):

What is the solution to the equation?\[-5 = \frac{ 8m + 12 }{ 4 }\] M = ???

OpenStudy (anonymous):

Alright. :3 Might only be able to answer one more because I have to go soon. xD

OpenStudy (anonymous):

B= 2/5 for number 2

OpenStudy (anonymous):

ok do this last one then :) and ill figure ou tthe rest :)

OpenStudy (anonymous):

M=-4

OpenStudy (anonymous):

What is the solution to the equation? \[-5 \frac{ 8m + 12 }{ 4 }\] M = ???

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Which answer best describes the solution to the equation 5z + 4 = 4 + 5z? A. one solution B. no solution C. an infinite number of solutions D. z = 20

OpenStudy (anonymous):

C an infinite number of solutions

OpenStudy (anonymous):

Which answer describes the solution to each equation? 1. \[\frac{ x - 6 }{ 2 } = x + 13\] 2. \[ 2 + 6x -7 = 8 + 2x - 10 + 4x\] 3. \[3x +5=\frac{ 1 }{ 5 }(15x+10)\] 4.\[\frac{ 1 }{ 4 }(16 + 8x) = 3x - 1 - x + 5\] A. one solution B. no solution C. an infinite number of solutions

OpenStudy (anonymous):

and yes it mean like 1 could be A and 2 could be B and 3 could be C and 4 could be B again

OpenStudy (anonymous):

do you understand it????

OpenStudy (anonymous):

ok im just gonna guess it :P i hope the rest of it was correct

OpenStudy (anonymous):

ok i mastered it so its good thanks though byee

OpenStudy (anonymous):

o ok sorry